User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!



Module: HOB+0121 0GB


Module: HOB+0083 0GB

  • approx 11:00 UT, antenna stuck. Resetting drives not working. Called Jamie who advised that it is likely due to weather. KE removed from dynamic observing at this stage. (JR)
  • If it is not up and working by ~23:00 UT, will contact Mick Stone and see if it is possible for him to make a trip out. (JR)
  • Attempted to contact Mick. Left a message. (JR)
  • 23:30 UT. Spoke to Warren who suggested that Brett may know what Mick needs to do to reset the drives. Brett to get into contact with Mick (who is currently on site) (JT)
  • 00:04 UT Issue resolved. Ke back is schedule. Fix: the power isolator to the antenna drives are located at knee height below the antenna drive cabinet in the pedestal. When the switch is vertical, the isolator is on. Mick switched it off momentarily (verified by 'it going quiet'), and then switched it back on. 10-20 seconds later HMI updated and the drives were OK again. Should you need to talk someone through this, get them to use the UTAS headset at yg (attached to x6283) (JR)
  • 08:11 UT - “wth” command doesn't return any values, a search for “wx” revealed it hasn't been working FOR THE ENTIRE EXPERIMENT - told Jamie. (Ellen)
  • 09:57 UT - Weather command now giving the 3 required values. Jamie fixed the wx reporting - the MET sensor webpage had locked up, so he changed the webpage address. (Ellen)
  • 10:37 UT - Delay difference dropped about 0.01us, but seems to be stable now. (Ellen)
  • 17:51 UT - alarm, fs self-terminated. Was restarted had to also restart ds and go schedule=ds351ke again (?). I think this worked first scan after crash will hopefully 352-1802 (Lucas).
  • There seems to have been a small clock jump just before 352/12:00UT (Bryn)


Module: HOB+0069 0GB

  • 00:45 UT alarm for no pcfs activity for 15 minutes. e-remote control lost connection and currently unable to vnc to pcfsyg (JR)
  • 01:05 UT connection re-established, e-control back up and able to ping timeyg and pcfsyg - dyshed script had to be restarted (JR)
  • 01:08 UT Back down again. (JR)
  • 01:17 UT reset VPN, still unable to connect
  • 01:30 UT rebooted D-Link router via vnc dyshed script restarted again, appears to be working (JR)
  • 04:28 UT (DOY353) - Yarragadee schedule script crashed, missed the current schedule file grab and is awaiting the new one. ( Tiege)
  • 05:43 UT - Turns out the problem I had initially thought was occurring was not the case, looks like pcfs could not write to the log anymore(?), potentially pcfs disk full, followed instructions outlined in the 'Current Issues' section of the wiki and deleted some files from /home/oper/auto_cor_test/ (Tiege)
  • 06:40 UT Yg finally back to normal operation. Tiege's problem at 05:43 UT somehow left the schedule=none. Hoped that would be fixed in the next dyn update interval. Wasn't. Re-started on pcfsyg. Schedule file recognised but a halt command was auto issued so yg remained idle. On the next dynamic update interval found source and all okay as of 06:40:20UT. (Liz)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/ds351.txt · Last modified: 2016/12/19 00:00 by Jonathan Rogers